OpenStudy (mew55):

integral 2 to infinity dx/x^2-1

myininaya (myininaya):

\[\int\limits_{2}^{\infty}\frac{dx}{x^2-1}\]?

myininaya (myininaya):

You can do partial fractions

OpenStudy (mew55):

oooh kk i can do this

OpenStudy (anonymous):

yeah, but since you will get the log is seems rather unlikely that it will converge, or am i missing something?

OpenStudy (anonymous):

-(log(x+1)-log(x-1))/2 , u do the rest

OpenStudy (anonymous):

oh yeah i am missing something let me shut up

OpenStudy (mew55):

XD

OpenStudy (mew55):

thank u all